A total of $8000 is invested: a part at 7% and remainder at 11%. How much is invested at each rate if the annual interest is $690?
Answer by mananth(16946)

Part I 7.00% per annum ------------- Amount invested =x
Part II 11.00% per annum ------------ Amount invested = y
8000
Interest----- 690.00

Part I 7.00% per annum ---x
Part II 11.00% per annum ---y
Total investment
x + 1 y= 8000 -------------1
Interest on both investments
7.00% x + 11.00% y= 690
Multiply by 100
7 x + 11 y= 69000.00 --------2
Multiply (1) by -7
we get
-7 x -7 y= -56000.00
Add this to (2)
0 x 4 y= 13000
divide by 4
y = 3250
Part I 7.00% $ 4750
Part II 11.00% $ 3250

CHECK
4750 --------- 7.00% ------- 332.50
3250 ------------- 11.00% ------- 357.50
Total -------------------- 690.00

m.ananth@hotmail.ca